# FAIntbench: A Holistic and Precise Benchmark for Bias Evaluation in Text-to-Image Models

### v1

@@ Abstract
Keywords: Text-to-Image Models, Bias Evaluation, Dataset, Distillation
@@ 1 Introduction
As one of the crucial multi-modal technologies in the field of AI-generated content (AIGC), Text-to-Image (T2I) generative models have attracted considerable interest (Rombach et al. (2022); Chen et al. (2023, 2024); Ding et al. (2022); Li et al. (2024); Luo et al. (2023a)). However, similar to the challenges encountered by large language models (Mehrabi et al. (2021); Gallegos et al. (2023)), biases in training datasets and algorithms also profoundly affect T2I models (Wan et al. (2024)). Research has indicated that even when supplied with prompts that lack specific protected attributes, T2I models primarily depict individuals with high social status occupations as white middle-aged men (Cho et al. (2023)).
Some researchers have conducted surveys or proposed their own solutions for decreasing biases (Bansal et al. (2022); Gandikota et al. (2024); Friedrich et al. (2023); Schramowski et al. (2023); Luccioni et al. (2023)). Obviously, researchers need a holistic bias benchmark to intuitively compare the biases of different models. However, existing benchmarks failed to fully meet the needs. For instance, DALL-EVAL by Cho et al. (2023) only evaluated biases towards occupations with 252 prompts, and HRS-Bench by Bakr et al. (2023) merely considered the situation that models failed to generate images with specific protected attribute. Additionally, the current benchmarks directly use the general definition of bias in the field of machine learning, lacking a definition and classification system particularly for T2I models.
@@
Table 2: The proportion distribution of the 2654 prompts in FAIntbench.
To address the issues outlined above, we introduce a holistic, precise, and adjustable bias benchmark named Fair AI Painter Benchmark, abbreviated as FAIntbench. Table 2 shows a simple comparison between existing benchmarks and our work. In FAIntbench, we establish a comprehensive definition system and classify biases across four dimensions. We construct the dataset with 2,654 prompts, covering occupations, characteristics and social relations, briefly depicted in Figure 2. FAIntbench employs fully automated evaluations based on the alignment by CLIP (Radford et al. (2021)), featuring adjustable evaluation metrics. The evaluation results cover implicit generative bias, explicit generative bias, ignorance, and discrimination. These characteristics make FAIntbench suitable for automated bias evaluation tasks for any T2I model. We evaluate seven recent large-scale T2I models with FAIntbench. Based on the results, we discuss the performance of the models in different biases and explore the effects of distillation (Meng et al. (2023)) on model biases. To ensure the reliability of the results of FAIntbench, we conduct human evaluations on a 10% sample, achieving significant consistency.
Our contributions are summarized as follows: (1) we establish a specific 4-dimension bias definition system for T2I models, which allows precise bias classification. (2) We introduce a dataset for comprehensive bias evaluation, which consists of a dataset including 2654 prompts. (3) We introduce FAIntbench, a holistic benchmark for bias evaluation in T2I models, including 18 evaluation metrics for biases covering 4 dimensions. (4) We conduct evaluations on seven models and human evaluation to prove the efficacy of FAIntbench. (5) We analyze the results to explore ongoing issues for biases in T2I models, for example, we verify that distillation increases biases.
@@ 2 FAIntbench Establishment
2.1 Definition System
To overcome the limitations of existing benchmarks that lack capability to classify and evaluate different biases (Bakr et al. (2023); Bansal et al. (2022)), we propose a new definition and classification system based on sociological and machine ethical studies on bias (Moule (2009); Landy et al. (2018); Varona and Suárez (2022); Chouldechova (2017)) and the guide provided by U.S. Department of Justice . Our definition system consists of four dimensions: manifestation of bias, visibility of bias, acquired attributes, and protected attributes. Detailed descriptions are provided in Appendix B.
Manifestation of Bias
@@ 2.1 Definition System
Visibility of Bias
From the perspective of the visibility of bias, we categorize bias into implicit generative bias and explicit generative bias. Their definition are inspired by implicit bias and explicit bias (Fridell (2013)) in sociology.
Acquired Attribute
An acquired attribute is a trait that individuals acquire through their experiences, actions, or choices.
Protected Attribute
A protected attribute is a shared identity of one social group, which are legally or ethically protected from being used as grounds for decision-making to prevent bias.
@@ 2.2 Dataset
We collect 2,654 prompts for our dataset. All of the prompts are generated by GPT-4 under supervising. We design our dataset along three dimensions. For the visibility dimension, we categorize prompts into two types: implicit prompts and explicit prompts. For the acquired attribute dimension, we include three attributes: occupation, social relation, and characteristic. For the protected attribute dimension, we also include three attributes: gender, race, and age. Figure 4 displays the complete construction formulas with corresponding examples.
Figure 1: Pipeline for the prompt set. The grey rectangle represents an identity prompt, the orange rectangle represents an photorealism prompt, the green rectangle represents the protected attributes, and yellow rectangle represents the acquired attributes.
To ensure the generated images suitable for evaluation, each of the 2,654 prompts consists of two sub-prompts: the identity prompt and the photorealism prompt. The identity prompt includes the identity of the individual depicted in each image and remains unique for each prompt. The photorealism prompt enhances the image’s realism. To accommodate compatibility, we exclude negative prompts. Additionally, we offer complete modification guidelines for customizing and updating the dataset. This adaptability enables FAIntbench to meet diverse research needs. Details of the dataset are available in Appendix C.
2.3 Evaluation Metrics
Our evaluation consists of three parts: alignment, implicit/explicit bias score evaluation, and manifestation factor evaluation. Figure 5 displays the pipeline of our evaluation. Our ground truth are selected from demographic statistics of U.S. Census Bureau (2022) and United Nations and Social Affairs (2022).
Figure 2: Pipeline for the evaluation. The yellow rectangle represents generated images, the black rectangle represents the meta data from alignment, the green box represents selected prompts for manifestation factor, and the red box represents the ground truth.
Alignment
In our alignment pipeline for generated images, each image is sequentially processed using optimized CLIP for alignment. First, we examine whether the images accurately depicted humans, setting a threshold to filter out and categorize failed images as hallucinations. Secondly, we use CLIP to align the images with protected attributes and use our optimization algorithm to improve the accuracy. Finally, we average the probability proportions across all images under each prompt to get the weights of protected attributes for this prompt. The pipeline of alignment is detailed in Appendix D.
@@ 2.3 Evaluation Metrics
Each prompt has an implicit bias score, ranging from 0 to 1, while higher scores indicate less bias. In the process of calculating the implicit bias score, we first retrieve the generative proportions of each protected attributes of the chosen prompt, alongside the corresponding demographic proportions of the prompt. We then calculate the cosine similarity between these sets of proportions and normalize it to produce the implicit bias score.
$S_{i,j}=\frac{1}{2}\left(\frac{\sum_{i=1}^{n}p_{i}\cdot q_{i}}{\sqrt{\sum_{i=1}^{n}p_{i}^{2}}\cdot\sqrt{\sum_{i=1}^{n}q_{i}^{2}}}+1\right)$
@@
$S_{sum,im}=\frac{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}k_{i}\cdot k_{j}\cdot S_{i,j}}{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}k_{i}\cdot k_{j}}$
@@
$S_{sum,ex}=\frac{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}k_{i}\cdot k_{j}\cdot S_{i,j}}{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}k_{i}\cdot k_{j}}$
@@
To evaluate whether biases of a model tend to ignorance or discrimination, we introduce a manifestation factor, denoted by $\eta$. Each protected attribute is assigned an $\eta$, with an initial value set to 0.5. The $\eta$ ranges from 0 to 1, as a lower $\eta$ indicates more ignorance while a higher $\eta$ suggests more discrimination. This initial value suggests that ignorance and discrimination contribute equally to the observed bias in the model.
@@
$\alpha_{i}=k_{i}\cdot((p_{i}-p^{\prime}_{i})^{2}+(q_{i}-q^{\prime}_{i})^{2})$
@@
$\eta=\eta_{0}+\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\begin{cases}\alpha_{i,j}&\text{if }((p_{i}>p^{\prime}_{i}\text{ and }q_{i}>q^{\prime}_{i})\text{ or }(p_{i}<p^{\prime}_{i}\text{ and }q_{i}<q^{\prime}_{i}))\\ -\alpha_{i,j}&\text{if }((p_{i}>p^{\prime}_{i}\text{ and }q_{i}<q^{\prime}_{i})\text{ or }(p_{i}<p^{\prime}_{i}\text{ and }q_{i}>q^{\prime}_{i}))\\ 0&\text{otherwise}\end{cases}$
@@ 3 Experiment
Setup
We evaluate the bias score of seven recent large-scale T2I models, i.e., SDXL (Podell et al. (2023)), SDXL-Turbo (Sauer et al. (2023)), SDXL-Lighting (Lin et al. (2024)), LCM-SDXL (Luo et al. (2023a)), PixArt-$\Sigma$ (Chen et al. (2024)), Playground V2.5 (Li et al. (2024)), and Stable Cascade (Pernias et al. (2023)). Each model is used to generate 800 images for each prompt to minimize the influence of chance on results. The parameters and additional results are shown in Appendix E.
@@ 3.1 Evaluation Results
Implicit Bias Result
Fig.6 and Fig.6 shows that SDXL has the best implicit bias score, while PixArt-$\Sigma$ performs the worst. The performance of SDXL-L, LCM-SDXL, Stable Cascade, and SDXL-T are close. For protected attributes, the performance of the seven models has similar traits, best in gender and worst in race, indicating a severe problem in racial biases.
@@
Explicit Bias Result
Fig.8 and Fig.8 shows that SDXL performs the best again, while PixArt-$\Sigma$ still performs the worst. The performance of SDXL-L, LCM-SDXL, and SDXL-T are close while Stable Cascade has the second highest score. For protected attributes, all models have the best performance in gender. SDXL-L and Pixart-$\Sigma$ perform wrost in race, while the others perform wrost in age.
@@
Manifestation Factor Result
The bias manifestations of all models tend to discrimination as Fig.10 depicted, suggesting that developers need to pay more attention to stereotypes to specific social groups.
3.2 Human Evaluation
To validate the reliability of FAIntbench, we select 265 prompts and 20 images for each prompt from the outputs of the SDXL and PixArt-$\Sigma$ for human evaluation. We have ten trained evaluators to align the protected attributes with the individuals in these images, and then we used the alignment results for bias calculations. The results validate the effectiveness of FAIntbench, as the average difference is 6.8%. Fig. 10 gives a clear comparison.
3.3 Discussion
@@
Additionally, It is important to note that due to different principles, implicit bias scores and explicit bias scores, although within the same range, can not be directly compared.

### v2

@@ 1 Introduction
As one of the crucial multi-modal technologies in the field of AI-generated content (AIGC), Text-to-Image (T2I) generative models have attracted considerable interest [43, 9, 8, 14, 29, 34]. However, similar to the challenges encountered by large language models [36, 19], biases in training datasets and algorithms also profoundly affect T2I models [52]. Research has indicated that even when supplied with prompts that lack specific protected attributes, T2I models primarily depict individuals with high social status occupations as white middle-aged men [11].
Some researchers have conducted surveys or proposed their own solutions for decreasing biases [4, 20, 18, 45, 33]. Obviously, researchers need a holistic bias benchmark to intuitively compare the biases of different models. However, existing benchmarks failed to fully meet the needs. For instance, DALL-EVAL by [11] only evaluated biases towards occupations with 252 prompts, and HRS-Bench by [3] merely considered the situation that models failed to generate images with specific protected attribute. Additionally, the current benchmarks directly use the general definition of bias in the field of machine learning, lacking a definition and classification system particularly for T2I models.
@@
To address the issues outlined above, we introduce a holistic, precise, and adjustable bias benchmark named Fair AI Painter Benchmark, abbreviated as FAIntbench. Table 1 shows a simple comparison between existing benchmarks and our work. In FAIntbench, we establish a comprehensive definition system and classify biases across four dimensions. We construct the dataset with 2,654 prompts, covering occupations, characteristics and social relations. FAIntbench employs fully automated evaluations based on the alignment by CLIP [41], featuring adjustable evaluation metrics. The evaluation results cover implicit generative bias, explicit generative bias, ignorance, and discrimination. These characteristics make FAIntbench suitable for automated bias evaluation tasks for any T2I model. We evaluate seven recent large-scale T2I models with FAIntbench. Based on the results, we discuss the performance of the models in different biases and explore the effects of distillation [37] on model biases. To ensure the reliability of the results of FAIntbench, we conduct human evaluations on a 10% sample, achieving significant consistency.
Our contributions are summarized as follows: (1) we establish a specific 4-dimension bias definition system for T2I models, which allows precise bias classification. (2) We introduce a dataset for comprehensive bias evaluation, which includes 2654 prompts. (3) We introduce FAIntbench, a holistic benchmark for bias evaluation in T2I models, covering 4 dimensions. (4) We conduct evaluations on seven models and human evaluation to prove the efficacy of FAIntbench. (5) We analyze the results to explore the side effect of distillation for biases in T2I models, i.e., verifying that distillation increases biases.
@@ 2 FAIntbench Establishment
2.1 Definition
To overcome the limitations of existing benchmarks that lack capability to classify and evaluate different biases [3, 4], we propose a new definition and classification system based on sociological and machine ethical studies on bias [38, 23, 25, 51, 12] and the guide provided by [50]. We consider our definition system from four dimensions: manifestation of bias, visibility of bias, acquired attributes, and protected attributes. Any kind of bias can be represented using these four dimensions.
2.1.1 Manifestation of Bias
@@
Ignorance refers to the phenomenon where T2I models consistently generate images depicting a specific demographic group, regardless of prompts suggesting positive and high-status terms or negative and low-status terms. This bias perpetuates a limited, homogenized view of diverse characteristics and roles, thereby reinforcing a narrowed societal perception.
Discrimination refers to the phenomenon where T2I models disproportionately associate positive and high-status terms with images of certain demographic groups, while aligning negative and low-status terms with images of other groups. This bias perpetuates and reinforces typical stereotypes to certain social groups.
2.1.2 Visibility of Bias
From the perspective of the visibility of bias, we categorize bias into implicit generative bias and explicit generative bias. Their definition are inspired by implicit bias [21] and explicit bias [17] in sociology.
Implicit generative bias refers to the phenomenon where, without specific instructions on protected attributes including gender, race, and age, T2I models tend to generate images that do not consist with the demographic realities. For instance, when a model is asked to generate images of a nurse, it only generate images of a female nurse without gender prompt.
Explicit generative bias refers to the phenomenon where, with specific instructions on protected attributes including gender, race, and age, T2I models tend to generate images that do not consist with the prompts. For instance, when a model is asked to generate images of an East-Asian husband with a white wife, it only generate images of an East-Asian couple.
2.1.3 Acquired Attribute
An acquired attribute is a trait that individuals acquire through their experiences, actions, or choices. It can be changed over time through personal effort, learning, experience, or other activities. They can be used as a reasonable basis for decision-making, but also possible to be related to bias and stereotype. Typical protected attributes include occupation, social relation, educational attainment, and personal wealth.
2.1.4 Protected Attribute
A protected attribute is a shared identity of one social group, which are legally or ethically protected from being used as grounds for decision-making to prevent bias. It is difficult to change as it is usually related to physiological traits. Typical protected attributes include race, gender, age, religion, and disability status.
@@ 2.2 Dataset
We collect 2,654 prompts for our dataset, briefly depicted in Figure 1. We design our dataset along three dimensions. For the visibility dimension, we categorize prompts into two types: implicit prompts and explicit prompts. For the acquired attribute dimension, we include three attributes: occupation, social relation, and characteristic. For the protected attribute dimension, we also include three attributes: gender, race, and age. More details are available in Appendix B.
Figure 1: The proportion distribution of the 2654 prompts in FAIntbench.
To ensure the generated images suitable for evaluation, each of the 2,654 prompts consists of two parts: the identity prompt and the photorealism prompt. The identity prompt includes the identity of the individual depicted in each image and remains unique for each prompt. The photorealism prompt enhances the image’s realism. To accommodate compatibility, we exclude negative prompts. Additionally, we offer complete modification guidelines for customizing and updating the dataset in the supplemental material. This adaptability enables FAIntbench to meet diverse research needs. Figure 2 displays the complete construction pipeline with corresponding examples.
Figure 2: Generation pipeline for the prompt set. Black pluses represent inserting attributes to identity prompts and blue plus represent the connection of prompts by GPT-4 [1].
2.3 Evaluation
Our evaluation consists of two parts: alignment and evaluation metrics. Figure 3 displays the pipeline of our evaluation. Our ground truth are selected from demographic statistics of [49] and [47].
Figure 3: Pipeline for the evaluation. The yellow rectangle represents generated images, the black box represents the meta data from alignment, the green box represents selected prompts for manifestation factor, and the red box represents the ground truth.
2.3.1 Alignment
In our alignment pipeline for generated images, each image is sequentially processed using optimized CLIP for alignment. First, we examine whether the images accurately depicted humans, setting a threshold to filter out and categorize failed images as hallucinations. Secondly, we use CLIP to align the images with protected attributes and use our optimization algorithm to improve the accuracy. Finally, we average the probability proportions across all images under each prompt to get the weights of protected attributes for this prompt. The pipeline of alignment is detailed in Appendix C.
2.3.2 Evaluation Metrics
Our evaluation metrics include three parts: implicit bias score evaluation, explicit bias score evaluation, and manifestation factor evaluation. Implicit and explicit bias scores reflect the severity of the biases in the models. They range from 0 to 1, while higher scores indicate less bias. The manifestation factor indicates whether biases of a model tend to ignorance or discrimination, denoted by $\eta$. The $\eta$ also ranges from 0 to 1, as a lower $\eta$ indicates more ignorance while a higher $\eta$ suggests more discrimination. We believe that these three metrics cover all common bias. We detail them in following sections.
@@ 2.3.2 Evaluation Metrics
In the process of calculating the implicit bias score, we first retrieve the generative proportions of each protected attributes of the chosen prompt, alongside the corresponding demographic proportions of the prompt. We then calculate the cosine similarity between these sets of proportions and normalize it to produce the implicit bias score.
$S_{i,j}=\frac{1}{2}\left(\frac{\sum_{i=1}^{n}p_{i}\cdot q_{i}}{\sqrt{\sum_{i=1}^{n}p_{i}^{2}}\cdot\sqrt{\sum_{i=1}^{n}q_{i}^{2}}}+1\right),$
@@
$S_{sum,im}=\frac{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}k_{i}\cdot k_{j}\cdot S_{i,j}}{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}k_{i}\cdot k_{j}},$
@@
$S_{sum,ex}=\frac{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}k_{i}\cdot k_{j}\cdot S_{i,j}}{\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}k_{i}\cdot k_{j}},$
@@
Each protected attribute is assigned an $\eta$, with an initial value set to 0.5. This initial value suggests that ignorance and discrimination contribute equally to the observed bias in the model.
@@
$\alpha_{i}=k_{i}\cdot((p_{i}-p^{\prime}_{i})^{2}+(q_{i}-q^{\prime}_{i})^{2}),$
@@
$\eta=\eta_{0}+\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\begin{cases}\alpha_{i,j}&\text{if }((p_{i}>p^{\prime}_{i}\text{ and }q_{i}>q^{\prime}_{i})\text{ or }(p_{i}<p^{\prime}_{i}\text{ and }q_{i}<q^{\prime}_{i}))\\ -\alpha_{i,j}&\text{if }((p_{i}>p^{\prime}_{i}\text{ and }q_{i}<q^{\prime}_{i})\text{ or }(p_{i}<p^{\prime}_{i}\text{ and }q_{i}>q^{\prime}_{i}))\\ 0&\text{otherwise}\end{cases},$
@@ 3 Experiment
We evaluate the bias score of seven recent large-scale T2I models, i.e., SDXL [40], SDXL-Turbo [44], SDXL-Lighting [31], LCM-SDXL [34], PixArt-$\Sigma$ [8], Playground V2.5 [29], and Stable Cascade [39]. Each model is used to generate 800 images for each prompt to minimize the influence of chance on results. The parameters and additional results are shown in Appendix D.
@@ 3.1 Evaluation Results
We briefly display our cumulative results in Table 2. These results indicate that the recent models perform well overall. However, there are certain differences among these models in specific metrics. We discuss these differences in following sections. Additionally, it is important to note that due to different principles, implicit bias scores and explicit bias scores, although within the same range, can not be directly compared.
Table 2: Qualitative cumulative results.
Attribute SDXL SDXL-L SDXL-T LCM PixArt Cascade PG2.5
Implicit Bias Score 0.875848 0.862829 0.867335 0.867246 0.849457 0.867106 0.848806
Explicit Bias Score 0.761633 0.717484 0.72936 0.703329 0.699121 0.754550 0.722061
Manifestation Factor 0.612138 0.620129 0.569934 0.624898 0.616241 0.617573 0.630034
3.1.1 Implicit Bias Result
Figure 7 and Figure 7 shows that SDXL has the best implicit bias score, while PixArt-$\Sigma$ performs the worst. The performance of SDXL-L, LCM-SDXL, Stable Cascade, and SDXL-T are close. For protected attributes, the performance of the seven models has similar traits, best in gender and worst in race, indicating a severe problem in racial biases.
@@ 3.1.1 Implicit Bias Result
We provide a typical instance for further illustration. When being requested to generate images of "an attractive person", all models tend to generate images of European and young people, while showing significant difference in gender. PixArt-$\Sigma$ has the best performance in gender while SDXL and its distillated models have similar performance. Stable Cascade almost ignores men. The quantitative results is shown in Table 3 and Figure 8.
Table 3: Qualitative results of the chosen implicit bias scores of the prompt "an attractive person".
Attribute SDXL SDXL-L SDXL-T LCM PixArt Cascade PG2.5
European 0.578283 0.868367 0.622451 0.670661 0.554337 0.739564 0.783
Woman 0.720386 0.669477 0.623881 0.737329 0.530357 0.949218 0.456293
Young 0.822438 0.80628 0.858108 0.786835 0.806469 0.817228 0.803891
Figure 8: Qualitative visualized results of the prompt "an attractive person".
3.1.2 Explicit Bias Result
Figure 10 and Figure 10 shows that SDXL performs the best again, while PixArt-$\Sigma$ still performs the worst. The performance of SDXL-L, LCM-SDXL, and SDXL-T are close while Stable Cascade has the second highest score. For protected attributes, all models have the best performance in gender. SDXL-L and Pixart-$\Sigma$ perform wrost in race, while the others perform wrost in age.
@@ 3.1.2 Explicit Bias Result
We provide a typical instance for further illustration with Figure 11. When being requested to generate images of "one East-Asian husband with one European wife", all models fail to generate correct images. Nevertheless, as a comparison, the models are mostly capable to correctly generate images of "one European husband with one East-Asian wife". This phenomena is consistent with a widespread stereotype, i.e., East-Asian men lack attractiveness to women of other ethnicities, which makes it difficult for them to find non-Asian spouses [28]. Recent research has shown that although the number of couples with Asian husbands and White wives is indeed lower than that of couples with White husbands and Asian wives, the difference is not such significant [32], which indicates that this stereotype represents a certain racial discrimination [27, 7]. However, all models exhibit significant bias on this issue, indicating the necessity for further research.
Figure 11: Qualitative visualized results of the prompts "one East-Asian husband with one European wife" and "one European husband with one East-Asian wife".
3.1.3 Manifestation Factor Result
The bias manifestations of all models tend to discrimination as Figure 13 depicted, suggesting that developers need to pay more attention to stereotypes to specific social groups. This result is consistent with our sampling estimation of the generated results.
3.2 Human Evaluation
To validate the reliability of FAIntbench, we select 265 prompts and 20 images for each prompt from the outputs of the SDXL and PixArt-$\Sigma$ for human evaluation. We have ten trained evaluators to align the protected attributes with the individuals in these images, and then we used the alignment results for bias calculations. The results validate the effectiveness of FAIntbench, as the average difference is 6.8%. Figure 13 gives a clear comparison.
@@ 3.2 Human Evaluation
3.3 Discussion of Distillation
@@
4 Limitations and Potential Improvements to FAIntbench
Based on the initial results, we propose the following potential future directions to delve deeper into the biases and their impacts within large-scale T2I models.
Test More Models: Due to budget constraints, models like DALL-E V2/V3 [42, 6], Midjourney v6, and Stable Diffusion V3 [15] have not been tested. Like all existing bias benchmarks, FAIntbench dosen’t evaluate models that are specifically optimized for debiasing by recent research [13, 20, 45]. Future studies will aim to include more models, especially those tailored to decrease biases.
Solve the Increasing Biases from Distillation: We have conducted preliminary studies and discussions on the side effects of distillation on bias using FAIntbench. Given the increasing focus on distillation techniques in the text-to-image domain ([46, 31, 35]), we believe that further research and tailored solutions for the increasing biases are crucial.